# From Images to Signals: Are Large Vision Models Useful for Time Series Analysis?

### v1

@@ 3 Methods for Using LVMs in Time Series Analysis
For TSF task, we employ two frameworks from the literature. Fig. 1(c) trains a linear forecaster zeng2023pixels ; yang2024vitime , Fig. 1(d) uses LVMs’ reconstruction decoders for forecasting chen2024visionts . Because only MAE and SimMIM in our study have such decoders, Fig. 1(d) is applied to them. Fig. 1(c) applies to ViT and Swin. For both frameworks, we adopt the “variate-independence” assumption that is widely used in TSF nie2023time , i.e., each variate is forecasted independently. This applies to all imaging methods except for MVH, by which all variates appear in the same image thus are forecasted once. Additionally, the framework in Fig. 1(d) adds a mask subsequent to the look-back window part in the image, then it reconstructs the masked patches and recovers forecasts. This requires input images to preserve raw time series values in pixels. Among the 8 imaging methods, only MVH and UVH preserve time series values. Thus, this framework is only applied to MVH and UVH. The framework in Fig. 1(c) can be applied to all imaging types. Table 1 summarizes how frameworks (b)(c)(d) in Fig. 1 apply to different LVMs.
@@ 4.3 In-Depth Analysis of LVMs’ Suitability in Time Series Tasks
[RQ6] What are the computational costs of LVMs? We evaluate the training and inference time of LVMs. Training time is measured when a model converges with early stopping. Inference time is estimated by the average runtime per test sample. Table 5 compares LVMs with the best two baselines in TSC (Fig. 2) – GPT4TS, TimesNet, and TSF (Table 2) – PatchTST, GPT4TS. From Table 5, LVMs have more parameters than the baselines. On average, LVMs take 6x (15x) training time than the best TSC (TSF) baseline, primarily due to their larger sizes of trainable parameters. For inference, LVMs are 0.5x faster than the best TSC baseline, but are 24x slower than the best TSF baseline. This is incurred by both the parameter size and the extra costs to imaging time series. Fig. 6 shows inference time vs. performance. Compared to the best baselines, LVMs trade the computational overhead for better performance. This is also evident in Fig. 2 and Table 2. Considering the fast developing hardware, the results suggest a big potential of LVMs in future time series research.
Figure 6: Inference time vs. performance of compared methods on TSC (accuracy) using Handwriting, SpokenArabicDigits, and TSF (MSE) using ETTh1, Weather. Full results are in Appendix B.10.
@@ B.4 Full Results of RQ2: Which imaging method best fits TSC (TSF) task?
This section provides detailed performance comparison of 8 imaging methods, including GAF, MVH, RP, STFT, Wavelet (Wave.), Filterbank (Filter.), UVH, and Line Plot. The best LVMs for TSC (i.e., ViT) and TSF (i.e., MAE) are used. Table 15 and Table 16 summarize the results for TSC and TSF, respectively. Unlike TSF, in which UVH demonstrates a clear advantage on the four datasets in Table 16, for TSC, we find ranking the compared methods using critical difference diagram (Fig. 4) over all TSC benchmark datasets gives higher confidence to identify the best imaging method (i.e., GAF). Thus Table 15 includes the results on all TSC benchmark datasets.
@@
Imaging Method GAF MVH RP STFT Wave. Filter. UVH Lineplot
ETTh1 96 0.986 0.783 0.484 0.471 0.969 0.771 0.534 0.533 0.621 0.582 0.820 0.684 0.356 0.383 0.902 0.751
192 1.004 0.797 0.575 0.517 0.971 0.775 0.621 0.587 0.650 0.600 0.864 0.707 0.395 0.406 1.204 0.894
336 1.038 0.820 0.623 0.546 0.989 0.788 0.602 0.573 0.681 0.616 0.827 0.693 0.417 0.424 1.223 0.901
720 1.008 0.812 0.737 0.612 1.062 0.825 0.669 0.621 0.699 0.633 0.858 0.720 0.467 0.463 1.150 0.852
Average 1.009 0.803 0.605 0.537 0.998 0.790 0.607 0.579 0.663 0.608 0.842 0.701 0.409 0.419 1.120 0.850
ETTm1 96 0.836 0.729 0.310 0.352 0.849 0.719 0.420 0.470 0.449 0.490 0.793 0.648 0.284 0.333 0.842 0.735
192 0.830 0.717 0.386 0.400 0.865 0.726 0.466 0.496 0.504 0.524 0.798 0.649 0.328 0.363 0.840 0.726
336 0.853 0.725 0.393 0.402 0.872 0.728 0.506 0.519 0.532 0.535 0.883 0.690 0.357 0.384 0.841 0.726
720 0.865 0.726 0.488 0.467 0.928 0.754 0.543 0.536 0.586 0.563 0.899 0.703 0.411 0.417 0.872 0.741
Average 0.846 0.724 0.394 0.405 0.879 0.732 0.484 0.505 0.518 0.528 0.843 0.673 0.345 0.374 0.849 0.732
Illness 24 5.066 1.591 2.326 0.976 5.106 1.594 5.049 1.591 4.270 1.484 7.863 2.056 1.977 0.921 4.993 1.508
36 5.236 1.628 2.152 0.919 5.309 1.629 5.143 1.598 4.293 1.487 8.169 2.122 1.812 0.872 5.147 1.593
48 5.118 1.600 2.111 0.966 5.381 1.643 5.010 1.574 4.190 1.451 7.144 1.962 1.743 0.856 5.039 1.541
60 5.349 1.641 2.118 0.968 5.586 1.685 5.164 1.601 4.045 1.430 7.193 1.986 1.816 0.881 5.235 1.601
Average 5.192 1.615 2.177 0.957 5.346 1.638 5.092 1.591 4.200 1.463 7.592 2.032 1.837 0.883 5.104 1.561
Weather 96 0.581 0.554 0.153 0.202 0.647 0.610 0.202 0.294 0.224 0.312 0.515 0.488 0.146 0.191 0.588 0.561
192 0.598 0.567 0.194 0.241 0.649 0.607 0.251 0.336 0.273 0.354 0.516 0.488 0.194 0.238 0.604 0.574
336 0.593 0.558 0.239 0.275 0.674 0.619 0.294 0.364 0.330 0.388 0.505 0.484 0.243 0.275 0.601 0.568
720 0.611 0.574 0.337 0.344 0.640 0.593 0.364 0.413 0.411 0.433 0.513 0.499 0.318 0.328 0.617 0.582
Average 0.596 0.563 0.231 0.266 0.653 0.607 0.278 0.352 0.310 0.372 0.512 0.490 0.225 0.258 0.603 0.571
Table 16: MSE and MAE comparison of 8 imaging methods on TSF benchmark datasets. Red numbers indicate the best performance for each dataset.
@@ B.7 Full Results of RQ5: Do LVMs capture temporal order of time series?
B.8 Full Results of RQ7: Which component of LVMs contributes more to forecasting
@@
B.9 Full Results of RQ8: Will period-based imaging method induce any bias?
@@
B.10 Full Results of RQ6: What are the computational costs of LVMs?

### v2

@@ 3 Methods for Using LVMs in Time Series Analysis
For TSF task, we employ two frameworks from the literature. Fig. 1(c) trains a linear forecaster zeng2023pixels ; yang2024vitime , Fig. 1(d) uses LVMs’ reconstruction decoders for forecasting chen2024visionts . Because only MAE and SimMIM in our study have such decoders, Fig. 1(d) is applied to them. Fig. 1(c) applies to ViT and Swin. For both frameworks, we adopt the “variate-independence” assumption that is widely used in TSF nie2023time , i.e., each variate is forecasted independently. This applies to all imaging methods except for MVH, by which all variates appear in the same image thus are forecasted once. Additionally, the framework in Fig. 1(d) adds a mask subsequent to the look-back window part in the image, then it reconstructs the masked patches and recovers forecasts. This requires input images to preserve raw time series values in pixels. Among the 8 imaging methods, only MVH and UVH preserve time series values. Thus, this framework is applied to MVH and UVH [^1] GAF can be applied as it has an inverse function, but is largely limited for reasons described in Appendix B.4.. The framework in Fig. 1(c) can be applied to all imaging types. Table 1 summarizes how frameworks (b)(c)(d) in Fig. 1 apply to different LVMs.
@@ 4.3 In-Depth Analysis of LVMs’ Suitability in Time Series Tasks
[RQ6] What are the computational costs of LVMs? We evaluate the training and inference time of LVMs. Training time is measured when a model converges with early stopping. Inference time is estimated by the average runtime per test sample. Table 5 compares LVMs with the best two baselines in TSC (Fig. 2) – GPT4TS, TimesNet, and TSF (Table 2) – PatchTST, GPT4TS. From Table 5, LVMs have more parameters than the baselines. On average, LVMs take 3x (16x) training time than the best TSC (TSF) baseline, primarily due to their larger sizes of trainable parameters. For inference, LVMs are 4x faster than the best TSC baseline, but are 20x slower than the best TSF baseline. This is incurred by both the parameter size and the extra costs to imaging time series. Fig. 6 shows inference time vs. performance. Compared to the best baselines in TSF, LVMs trade the computational overhead for better performance. This is also evident in Fig. 2 and Table 2. Considering the fast developing hardware, the results suggest a big potential of LVMs in future time series research.
Figure 6: Inference time vs. performance of compared methods on TSC (accuracy) using UWaveGesture, SpokenArabicDigits, and TSF (MSE) using ETTh1, Weather. Full results are in Appendix B.8.
@@ B.4 Full Results of RQ2: Which imaging method best fits TSC (TSF) task?
This section provides detailed performance comparison of 8 imaging methods, including GAF, MVH, RP, STFT, Wavelet (Wave.), Filterbank (Filter.), UVH, and Line Plot. The best LVMs for TSC (i.e., ViT) and TSF (i.e., MAE) are used. Table 15 and Table 16 summarize the results for TSC and TSF, respectively. For TSF, UVH demonstrates a clear advantage on the 4 datasets in Table 16. For TSC, all benchmark datasets are used in Table 15 because the 4 datasets outlined in §$\S$4.3 are insufficient to confidently rank the compared methods using critical difference (CD) diagram (Fig. 4). Using all datasets improves confidence and helps identify the best imaging method (i.e., GAF).
In Table 16, GAF (GAF) represents applying GAF with framework (c) (framework (d)) in Fig. 1. For GAF, we follow [48] to use its inverse function to recover forecasted values from the reconstructed images by the framework in Fig. 1(d). Notably, GAF scales all time series values within [0, 1] by min-max normalization to compute polar coordinates during its imaging process. The normalization uses the minimum and maximum values from the look-back window, which are used to recover any predicted values. This imposes a constraint on the predicted values, i.e., the predicted values must remain within the upper and lower bounds of the look-back window, which is irrational in TSF, leading to a significant limitation and performance degradation as demonstrated in Table 16. As such, we use GAF, which outperforms GAF in Table 16, in the CD diagram in Fig. 4(b).
@@
Imaging Method GAF GAF MVH RP STFT Wave. Filter. UVH Lineplot
Dataset Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.986 0.783 1.224 0.850 0.484 0.471 0.969 0.771 0.534 0.533 0.621 0.582 0.820 0.684 0.356 0.383 0.902 0.751
192 1.004 0.797 1.227 0.854 0.575 0.517 0.971 0.775 0.621 0.587 0.650 0.600 0.864 0.707 0.395 0.406 1.204 0.894
336 1.038 0.820 1.214 0.857 0.623 0.546 0.989 0.788 0.602 0.573 0.681 0.616 0.827 0.693 0.417 0.424 1.223 0.901
720 1.008 0.812 1.190 0.863 0.737 0.612 1.062 0.825 0.669 0.621 0.699 0.633 0.858 0.720 0.467 0.463 1.150 0.852
Average 1.009 0.803 1.214 0.856 0.605 0.537 0.998 0.790 0.607 0.579 0.663 0.608 0.842 0.701 0.409 0.419 1.120 0.850
ETTm1 96 0.836 0.729 0.956 0.676 0.310 0.352 0.849 0.719 0.420 0.470 0.449 0.490 0.793 0.648 0.284 0.333 0.842 0.735
192 0.830 0.717 0.967 0.685 0.386 0.400 0.865 0.726 0.466 0.496 0.504 0.524 0.798 0.649 0.328 0.363 0.840 0.726
336 0.853 0.725 0.988 0.697 0.393 0.402 0.872 0.728 0.506 0.519 0.532 0.535 0.883 0.690 0.357 0.384 0.841 0.726
720 0.865 0.726 1.107 0.779 0.488 0.467 0.928 0.754 0.543 0.536 0.586 0.563 0.899 0.703 0.411 0.417 0.872 0.741
Average 0.846 0.724 1.005 0.709 0.394 0.405 0.879 0.732 0.484 0.505 0.518 0.528 0.843 0.673 0.345 0.374 0.849 0.732
Illness 24 5.066 1.591 6.172 2.618 2.326 0.976 5.106 1.594 5.049 1.591 4.270 1.484 7.863 2.056 1.977 0.921 4.993 1.508
36 5.236 1.628 5.497 2.627 2.152 0.919 5.309 1.629 5.143 1.598 4.293 1.487 8.169 2.122 1.812 0.872 5.147 1.593
48 5.118 1.600 5.218 2.448 2.111 0.966 5.381 1.643 5.010 1.574 4.190 1.451 7.144 1.962 1.743 0.856 5.039 1.541
60 5.349 1.641 5.299 2.239 2.118 0.968 5.586 1.685 5.164 1.601 4.045 1.430 7.193 1.986 1.816 0.881 5.235 1.601
Average 5.192 1.615 5.547 2.483 2.177 0.957 5.346 1.638 5.092 1.591 4.200 1.463 7.592 2.032 1.837 0.883 5.104 1.561
Weather 96 0.581 0.554 0.961 0.592 0.153 0.202 0.647 0.610 0.202 0.294 0.224 0.312 0.515 0.488 0.146 0.191 0.588 0.561
192 0.598 0.567 0.995 0.614 0.194 0.241 0.649 0.607 0.251 0.336 0.273 0.354 0.516 0.488 0.194 0.238 0.604 0.574
336 0.593 0.558 1.039 0.637 0.239 0.275 0.674 0.619 0.294 0.364 0.330 0.388 0.505 0.484 0.243 0.275 0.601 0.568
720 0.611 0.574 1.051 0.644 0.337 0.344 0.640 0.593 0.364 0.413 0.411 0.433 0.513 0.499 0.318 0.328 0.617 0.582
Average 0.596 0.563 1.012 0.622 0.231 0.266 0.653 0.607 0.278 0.352 0.310 0.372 0.512 0.490 0.225 0.258 0.603 0.571
Table 16: MSE and MAE comparison of 8 imaging methods on TSF benchmark datasets. Red numbers indicate the best performance for each dataset. GAF represents applying GAF with the framework in Fig. 1(c). GAF represents applying GAF with the framework in Fig. 1(d).
@@ B.7 Full Results of RQ5: Do LVMs capture temporal order of time series?
B.8 Full Results of RQ6: What are the computational costs of LVMs?
B.9 Full Results of RQ7: Which component of LVMs contributes more to forecasting
@@
B.10 Full Results of RQ8: Will period-based imaging method induce any bias?